OpenStudy (anonymous):

find the 23th term of the arithmetic sequence -1,-3,-5,-7

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

a=-1 d=-2 nth term=a +(n-1)d 23rd term=-1+(23-1)-2=-1-44=-45

OpenStudy (anonymous):

thank you , but i have another question also

OpenStudy (anonymous):

find the 20th term of the airthmatic sequence 2.5, 5, 7.5, 10 , 12.5

OpenStudy (anonymous):

a=2.5 d=2.5 20th term =2.5+19x2.5=50

OpenStudy (anonymous):

IF THE FIRST TERM OF AN ARITHMATIC SEQUENCE IS 6, AND 22, WHAT IS THE COMMON DIFFRENCE

OpenStudy (anonymous):

umm 22 wat is 22??

OpenStudy (anonymous):

THE FIFTH TERM

OpenStudy (anonymous):

well then 22=a +4d,22=6+4d,16=4d,d=4

OpenStudy (anonymous):

THANKS , If the first term of an arithmetic sequence is 10 and the 13th term is 58, what is the common difference?

OpenStudy (anonymous):

umm i guess u shud start pluggin d values in2 d equation a=(n-1)d inorder 2 get a hang at such q's!however d ans is 58=10+12d,d=4

OpenStudy (anonymous):

Describe the boundary curve and shading for the inequality. y < x2 - 12x + 20 a) boundary solid; shading outside b) boundary solid; shading inside c) boundary dashed; shading outside d) boundary dashed; shading inside

OpenStudy (anonymous):

umm srry bt i m pretty clueless abt boundary curves!!

OpenStudy (anonymous):

WHAT ABOUT FINDING THE VERTEX OF A PARABOLA?

OpenStudy (anonymous):

Find the vertex and tell whether the graph opens up or down. y > x2 - 2x - 3 a) (-1, -4); opens down b) (-1, -4); opens up c) (1, -4); opens down d) (1, -4); opens up

OpenStudy (anonymous):

DO YOU THINK YOU KNOW HOW TO DO THAT AT ALL ANY ???

OpenStudy (anonymous):

HELLO??

OpenStudy (ajprincess):

May b this may b of some use to u.

OpenStudy (anonymous):

@ajprincess thank you so much god bless you ! :)

OpenStudy (ajprincess):

yw nd thanx.